The differential equation satisfied by a beam uniformly loaded at one end that is fixed and the other end 8s subjected to a tensile force P, is given by EI.d^2y/dx^2= P.y- 1/2(W.x^2),where E,I,P,W are constants. Show that the elastic curve for the beam under the conditions y=0, dy/dx=0 at x=0, is given by y(x)=(W/P^2)[1-cosh(vx)]+ (W/2P)(x^2+2/n^2),where EI= P/n^2
1
Expert's answer
2020-11-13T15:55:48-0500

"\displaystyle
\textsf{Consider the differential equation}\\

ay" = py- bx^2\\hspace{1cm}(1)\\\\\n\nay" - py = -bx^2\\

\textsf{The auxiliary equation is}\\
ak^2 - p = 0\\

k^2 = \frac{p}{a}\\

k = \pm\sqrt{\frac{p}{a}}\\

\textsf{Since}\,\, p, a > 0\\

y = CF = A\cosh\left(\sqrt{\frac{p}{a}}x\right) + B\sinh\left(\sqrt{\frac{p}{a}}x\right)\\

ay" - py = -bx^2\\\\\n\n\ny = PI = lx^2 + mx + n\\\\\n\ny" = 2l\\


\implies 2la - p(lx^2 + mx + n) = -bx^2\\


pl = b, l = \frac{b}{p}\\

pm = 0, m = 0\\

2la - pn = 0, \\

n = \frac{2la}{p} = \frac{2ab}{p^2}\\


\therefore y = PI = \frac{b}{p}x^2 + \frac{2ab}{p^2}\\

y = CF + PI\\


\therefore y = A\cosh\left(\sqrt{\frac{p}{a}}x\right) + B\sinh\left(\sqrt{\frac{p}{a}}x\right) + \frac{b}{p}x^2 + \frac{2ab}{p^2}\\


0 = A + \frac{2ab}{p^2}\\

A = -\frac{2ab}{p^2}\\

y' = A\sqrt{\frac{p}{a}}\sinh\left(\sqrt{\frac{p}{a}}x\right) + B\sqrt{\frac{p}{a}}\cosh\left(\sqrt{\frac{p}{a}}x\right) + \frac{2bx}{p}\\

y'(0) = 0 = B\sqrt{\frac{p}{a}}\\

\therefore B = 0\\


y = -\frac{2ab}{p^2}\cosh\left(\sqrt{\frac{p}{a}}x\right)+ \frac{b}{p}x^2 + \frac{2ab}{p^2} \\

y = \frac{2ab}{p^2}\left(1 - \cosh\left(\sqrt{\frac{p}{a}}x\right)\right) + \frac{b}{p}x^2 \\

\textsf{Comparing the DE in}\, (1)\, \, \textsf{with the given DE.}\\

ay" = py- bx^2\\\\\ny"= P.y- 1/2(W.x^2)\\


b = \frac{W}{2}, p = P, a = El, v = \sqrt{\frac{P}{a}}\\

y = \frac{2El.W}{2P^2}\left(1 - \cosh(vx)\right) + \frac{W}{2P}x^2 \\

\therefore y = \frac{El.W}{P^2}\left(1 - \cosh(vx)\right) + \frac{W}{2P}x^2\\"